OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2? y = 2x + 1 y = 2x + 3 y = 2x - 8 y = 2x - 3

OpenStudy (anonymous):

y = mx + c, you know m =2 y = 2x + c you know it passes through (2,5) 5 = 2(2) + c, c = 1 therefore y = 2x + 1

OpenStudy (anonymous):

okay thanks